# A Joint X-ray and Optical Study of the Massive Redback Pulsar J2215+5135

### v1

@@ Abstract
PSR J2215+5135 (J2215) is a ‘redback’ spider pulsar, where the intrabinary shock (IBS) wraps around the pulsar rather than the stellar-mass companion. Spider orbital light curves are modulated, dominated by their binary companion thermal emission in the optical band and by IBS synchrotron emission in the X-rays. We report on new XMM-Newton X-ray and U-band observations of J2215. We produce orbital light curves and use them to model the system properties. From our optical light curve modeling, we find a neutron star mass $M_{NS}=2.15\pm 0.10M_{\odot}$, somewhat lower than previously reported. From the X-ray analysis, we find that the IBS still wraps around the pulsar, but with a wind/companion wind momenta ratio unusually close to unity, implying a flatter IBS geometry than seen in other spiders. Estimating the companion wind momentum and speed from the X-ray light curve, we find a companion mass-loss rate of ${\dot{M}}_{c}\gtrsim 10^{-10}$ M yr-1, so that J2215 may become an isolated millisecond pulsar in $\sim 1$ Gyr. Our X-ray analyses support models of magnetic reconnection and particle acceleration in the highly magnetized relativistic IBS.
@@ 1 Introduction
X-ray observations may probe the IBS while optical observations typically reveal the companion heating. Particles in the shocked pulsar wind accelerate and emit synchrotron X-rays (Kandel et al., 2019, 2021). The light curves are characterized by two caustic peaks per orbital phase, associated with the beamed emission from relativistic particles traveling tangent to the observing line of sight. Thermal emission from the companion is generally boosted by re-processed pulsar gamma-rays on the ‘day’ side of the companion and sometimes by precipitating IBS particles reprocessed in surface hot spots. Consequently, X-ray light curve and spectral analyses reveal the IBS structure and particle acceleration mechanisms, while optical analyses examine heating of the companion surface.
PSR J2215+5135 (J2215 hereafter) is a redback millisecond pulsar (MSP) with spin period $P_{s}=2.61$ s and spin-down power $\dot{E}=5\times 10^{34}$ $I_{45}$ erg s-1 in a $P_{b}=4.14$ hr orbit with a $M\approx 0.3$ M companion (Hessels et al., 2011). The source has dispersion measure of 69.2 pc cm-3 (Hessels et al., 2011), which corresponds to an estimated distance $\sim 3$ kpc (Yao et al., 2017, e.g.). First detected as an unidentified gamma-ray source by the Fermi Large Area Telescope (Fermi-LAT) (Atwood et al., 2009) and discovered by 350 MHz Green Bank Telescope followup observations (Ray et al., 2012; Hessels et al., 2011), J2215 has been studied optically (Breton et al., 2013; Schroeder & Halpern, 2014; Romani et al., 2015; Linares et al., 2018), in radio (Hessels et al., 2011; Broderick et al., 2016), in gamma-rays (Romani et al., 2015), and briefly in X-rays (Gentile et al., 2014; Linares, 2014) prior to this work.
Optical light curve modeling suggests that J2215 contains a particularly massive neutron star $\gtrsim 2$ M (Linares et al., 2018; Kandel & Romani, 2020). Kandel & Romani (2020) estimate a neutron star mass of $M_{NS}=2.24\pm 0.09$ M using a companion surface optical model that contains a magnetic pole hot spot. Such a massive neutron star constrains properties of the dense matter equation of state (Lattimer & Prakash, 2007; Steiner et al., 2013; Brandes et al., 2023). Previous modeling is sensitive to the poorly constrained companion hot spot, adding uncertainty to the binary inclination $i$ and, by extension, $M_{NS}$ estimates. Also, previous X-ray observations, which suggested widely separated light curve peaks and a flat IBS, relied on an XMM-Newton exposure with strong background flaring and only two binary orbits. These factors motivate further optical and X-ray study.
@@ 2.1 X-rays
We convert the PN and MOS light curves from count-rate and count-rate error to flux in each detector. We combine the data and distribute it into 26 phase bins. We compute the average flux weighted by the inverse square of the errors in each bin. Flux errors are estimated by the standard error of a variance-weighted sample. Our combined 0.5-10 keV X-ray light curve is shown in the top panel of Fig. 1.
Figure 1: Binned X-ray and optical orbital light curves of J2215. Less-finely binned light curves are shown by faded markers. (Top) The combined PN and MOS 0.5-10 keV X-ray light curve. We identify four distinct regions for which we separately fit spectra: two distinct peaks, P1 and P2 corresponding to the two phase intervals in which the flow of the IBS shock is tangent to our line of sight, a bridge (B) between the two peaks, and an off region around the companion inferior conjunction where the X-ray emission comes principally from the IBS nose. (Bottom) The OM U filter light curve for J2215 from the new XMM-Newton observations.
@@ 2.2 Optical
Contemporaneous OM U filter data was taken in both image and fast mode. The fast mode 10”$\times$10” images were obtained from a series of exposures over each observation. We use the SAS tool omfchain to obtain U filter light curves. For each exposure, we extract the source from a circular aperture with radius of 3 detector pixels in the fast mode image and determine the background from the average count rate of the corresponding image mode image with all point sources removed. After processing, the OM exposure time is 91 ks. We convert the OM U data to flux using the published OM U calibration for A-type stars. We place the data into flux bins by computing the variance weighted sample mean and standard error. We also use the Fermi-LAT third gamma-ray pulsar catalog ephemeris for the orbital parameters. The OM U light curve is shown in the bottom panel of Fig. 1.
@@ 3 Spectral Analysis
We divide our X-ray light curve into four phase intervals as labeled in Fig. 1 and perform spectral analysis of these phase bins. We identify Peak 1 (P1; $0.580.44<\Phi<$ 0.58) and Peak 2 (P2; $1.070.87<\Phi<$ 1.07) as the two characteristic IBS X-ray peaks (Romani & Sanchez, 2016; Kandel et al., 2019). We also demarcate a bridge region (B; $0.870.58<\Phi<$ 0.87) between P1 and P2 and an off region ($0.440.07<\Phi<$ 0.44) corresponding to the companion at inferior conjunction. Here non-IBS fluxes (e.g. from the magnetosphere or a pulsar wind nebula) may dominate, as only weakly beamed X-rays from the IBS nose are expected. We use the CIAO tool Sherpa (Freeman et al., 2001; Doe et al., 2007) to fit non-thermal power-law spectral models to the XMM data with spectral indices $\Gamma_{P1}$, $\Gamma_{P2}$, and $\Gamma_{B}$. We include an additional phase-independent power-law component throughout the orbit with power-law index $\Gamma_{0}$. The spectral model in counts s-1 keV-1 in each phase region is
@@
where $N_{H}$ is the equivalent hydrogen column, $\sigma(E)$ is the photo-electric cross section, $K$ is the power-law normalization factor, and $r=\{P1,P2,B\}$ demarcates the orbital phase region. For our absorption model, we employ the photoelectric cross sections of Balucinska-Church & McCammon (1992) and the elemental abundances of Anders & Grevesse (1989). In the off phase, $K_{r}=0$. We include the underlying power-law to account for phase-independent emission. The results of our fit are shown in Table 1. Because the spectral indices of P1, P2, and B are not significantly different, we also give the average spectral index across the IBS $\Gamma_{IBS}$. Fig. 2 shows the spectra across the IBS as well as in the off phases as observed by PN.
Figure 2: The observed PN X-ray spectra with the equivalent hydrogen column $N_{H}$ fixed at $0.09\times 10^{22}$ cm-2. The plotted points are the spectra in the IBS phase region ($1.070.44<\Phi<$ 1.07; including P1, P2, and B) as well as in the off region ($0.440.07<\Phi<$ 0.44). The solid line shows the full model in the IBS phases $f_{tot}$ and the two dashed lines show the absorbed power-law models for the IBS component $f_{IBS}$ with power-law index $\Gamma_{IBS}$ and the phase-independent component $f_{0}$ with power-law index $\Gamma_{0}$. Note that in the off region, the only model component is the absorbed power-law $f_{0}$.
@@ 4.1 Optical Models
Previous optical modeling of J2215 has been conducted by Linares et al. (2018) and Kandel & Romani (2020) using g’r’i’ data from two observing nights at the William Herschel Telescope (WHT). These light curve analyses show strong heating asymmetries on the companion star, suggesting the presence of a hot spot, which could plausibly be a magnetic pole (Sanchez & Romani, 2017). Kandel & Romani (2020) include a hot spot at the coordinates $(\theta_{HS},\phi_{HS})$ defined from the nose on the stellar surface. At that location, the stellar temperature increases by a factor $(1+A_{HS})$ from $T\approx 5600$ K to $T\approx 8000$ K with a Gaussian profile over radius $r_{HS}$. Their best-fit model gives $i=68.9^{\circ}$ and consequently a very high $M_{NS}=2.24$ M. These results, however, are sensitive to the hot spot parameters. Additionally, they were obtained using a simplified limb-darkening model, which does not account for the limb-darkening of the reprocessed pulsar heating luminosity (Claret & Bloemen, 2011; Kandel & Romani, 2023).
Figure 3: The best-fit companion optical light curve model plotted against the gri data from WHT and the OM U data shifted by the estimated calibration offsets along with the residuals. Phase $0\leq\Phi\leq 1$ shows WHT night 1 while $1\leq\Phi\leq 2$ shows WHT night 2.
We refit the WHT g’r’i’ data along with the new OM U data using the ICARUS light curve modeling code (Breton et al., 2013) including the extensions described in Kandel & Romani (2020) and Kandel & Romani (2023). We fit the light curves in each band simultaneously, using the same Fermi-LAT ephemeris since it is valid for the WHT data taken 2014 August 11 and 2014 September 1. We perform Monte Carlo sampling of the parameters using the library PyMultinest (Feroz & Hobson, 2008; Feroz et al., 2009; Buchner et al., 2014; Feroz et al., 2019), which uses nested sampling to search the parameter space.
@@
The model is very similar to the best-fit hot spot model of Kandel & Romani (2020); we augment a direct pulsar heating model with a surface hot spot, as well as add a flat spectrum veiling flux $f_{2}$ to the WHT night 2 data. This veiling flux was required by Kandel & Romani (2020), and is retained in our fits. Our fits prefer calibration offsets $\delta m_{r}\approx 0.03$ to both nights in the WHT r’ band and $\delta m_{U}\approx 0.66$ to the OM U data. We have attempted to find the origin of this large U offset. As noted in Sec. 2.2, we obtain the OM U AB magnitude by converting from count rate to flux with the SAS-recommended U conversion factor of $1.7\times 10^{-16}$ erg cm-2 Å-1 count-1 for a spectral type A stars. This spectral type is appropriate for the companion near optical maximum. Using archival XMM OM U exposures in fields covered by the Sloan Digital Sky Survey (SDSS), where catalog $u$ magnitudes are available, we have extracted stellar count rates following the same procedure used for J2215. From these data we find a conversion $2.7\pm 0.2\times 10^{-16}$ erg cm-2 Å-1 count-1, which would suggest that the companion U fluxes are brighter by a factor of $\sim 1.6$ (or offset by $\sim 0.5$ mag) with respect to the our XMM calibrated flux, close to the offset needed.
@@
We have also examined the calibrated 2014 Keck LRIS spectra of Romani et al. (2015), whose gri colors are consistent with the 2014 WHT photometry. We find the Keck fluxes at wavelengths below $4000$ consistent with BT-Settl model atmosphere spectra (Allard, 2014) at the day side temperature of the star. A change to the thermal companion flux (unlike the IBS non-thermal flux) seems unlikely and would require a concomitant change in the heating power. Our X-ray flux is consistent with past values and there is no significant change to the (dominant) Fermi-LAT gamma-ray heating. Consequently, we consider a U calibration offset the most likely explanation. While fitting for $\delta m_{U}$ slightly decreases the predictive power of our model, the OM U light curve shape nevertheless constrains the hot spot geometry and allows a good model fit. Of course, precisely calibrated fluxes and simultaneous UV/optical photometry would provide more secure constraints.
The fit-estimated parameters are shown in Table 2, while the best-fit model light curves are shown in Fig. 3. The $\chi^{2}/$DoF=1.22, while not perfect, is lower than 1.28 as reported by Kandel & Romani (2020). While the overall convergence of the fit has improved, our constraints on $i$ are weaker. This is a consequence of the different limb-darkening prescription used. Allowing the heated luminosity to be reprocessed in addition to the internal stellar flux tends to increase the flux on the night side of the star, thus requiring larger $i$ values to reproduce the night-day flux difference. The constraints on the hot spot parameters, however, have improved by a factor of 2, a result of both the different limb-darkening prescription and the additional U band data.
The larger value of $i$ implies a slightly lower neutron star mass than found by (Kandel & Romani, 2020). To determine the neutron star mass (also shown in Table 2) we use the Kandel & Romani (2020) radial velocity measurement $K_{c}=429.8\pm 3.9$ km/s. Our new mass estimate $M_{NS}=2.15\pm 0.10$ M is within a standard deviation of $M_{NS}=2.24\pm 0.09$ M. Consequently, J2215 likely remains $>2$ M and is thus useful for constraining the neutron star equation of state.
Parameters Results KR20 Results
@@
$\delta m_{r}$ $0.027^{+0.001}_{-0.001}$ $0.031^{+0.001}_{-0.001}$
$\delta m_{U}^{\dagger}$ $0.664^{+0.035}_{-0.036}$
$\theta_{HS}$ (deg) $340.5^{+6.0}_{-5.0}$ $324.0^{+10.1}_{-8.7}$
$\phi_{HS}$ (deg) $71.6^{+5.0}_{-10.0}$ $73.2^{+5.0}_{-8.4}$
$A_{HS}$ $0.35^{+0.16}_{-0.09}$ $0.6^{+0.4}_{-0.2}$
$r_{HS}$ (deg) $19.1^{+3.3}_{-6.4}$ $15.5^{+3.9}_{-6.8}$
$\chi^{2}$/DoF 311/255 297/232
Table 2: Parameter results for optical modeling of WHT+OM J2215 data including the 68% confidence intervals. $f_{c}$ is the Roche lobe filling factor, $L_{H}$ is the pulsar heating luminosity assuming an equatorially concentrated heating flux, and $d$ is the source distance. The hot spot coordinates $(\theta_{HS},\phi_{HS})$ are defined so that $(0,0)$ refers to the sub-pulsar point at the companion nose. The dust extinction value $A_{V}=0.40$ (Green et al., 2018) is used. The right column shows the fit results of Kandel & Romani (2020) (KR20) which use a different gravity darkening model and do not include U data. $\delta m_{U}$ is an offset from the standard SAS flux calibration; however, we see evidence for a systematic calibration shift of at least 0.5 mag.

### v2

# The Intrabinary Shock and Companion Star of Redback Pulsar J2215+5135

@@ Abstract
PSR J2215+5135 (J2215) is a ‘redback’ spider pulsar, where the intrabinary shock (IBS) wraps around the pulsar rather than the stellar-mass companion. Spider orbital light curves are modulated, dominated by their binary companion thermal emission in the optical bands and by IBS synchrotron emission in the X-rays. We report on new XMM-Newton X-ray and U-band observations of J2215. We produce orbital light curves and use them to model the system properties. Our best-fit optical light model gives a neutron star mass $M_{NS}=1.98\pm 0.08$ M, lower than previously reported. However, uncertainty in the stellar atmosphere metallicity, a parameter to which J2215 is unusually sensitive, requires us to consider an acceptable systematic plus statistical range of $M_{NS}\sim 1.85-2.3$ M. From the X-ray analysis, we find that the IBS wraps around the pulsar, but with a pulsar wind to companion wind momentum ratio unusually close to unity, implying a flatter IBS geometry than seen in other spiders. Estimating the companion wind momentum and speed from the X-ray light curve, we find a companion mass-loss rate of ${\dot{M}}_{c}\gtrsim 10^{-10}$ M yr-1, so that J2215 may become an isolated millisecond pulsar in $\sim 1$ Gyr. Our X-ray analyses place constraints on the magnetization and particle density of the pulsar wind and support models of magnetic reconnection and particle acceleration in the highly magnetized relativistic IBS.
@@ 1 Introduction
X-ray observations may probe the IBS while optical observations typically reveal the companion heating. Particles in the shocked pulsar wind accelerate and emit synchrotron X-rays (Kandel et al., 2019, 2021). The light curves are characterized by two caustic peaks per orbital phase, associated with the beamed emission from relativistic particles traveling tangent to the observing line of sight. Thermal emission from the companion is generally boosted by reprocessed pulsar gamma-rays on the ‘day’ side of the companion and sometimes by precipitating IBS particles reprocessed in surface hot spots. Consequently, X-ray light curve and spectral analyses reveal the IBS structure and particle acceleration mechanisms, while optical analyses examine heating of the companion surface.
PSR J2215+5135 (J2215 hereafter) is a redback millisecond pulsar (MSP) with spin period $P_{s}=2.61$ ms and spin-down power $\dot{E}=5\times 10^{34}$ $I_{45}$ erg s-1 in a $P_{b}=4.14$ hr orbit with a $M\approx 0.3$ M companion (Hessels et al., 2011). The source has dispersion measure of 69.2 pc cm-3 (Hessels et al., 2011), which corresponds to an estimated distance $\sim 3$ kpc (Yao et al., 2017, e.g.). First detected as an unidentified gamma-ray source by the Fermi Large Area Telescope (Fermi-LAT) (Atwood et al., 2009) and discovered by 350 MHz Green Bank Telescope followup observations (Ray et al., 2012; Hessels et al., 2011), J2215 has been studied optically (Breton et al., 2013; Schroeder & Halpern, 2014; Romani et al., 2015; Linares et al., 2018), in radio (Hessels et al., 2011; Broderick et al., 2016), in gamma-rays (Romani et al., 2015), and briefly in X-rays (Gentile et al., 2014; Linares, 2014) prior to this work.
Earlier optical light curve modeling suggests that J2215 contains a particularly massive neutron star $\gtrsim 2$ M (Linares et al., 2018; Kandel & Romani, 2020, hereafter KR20). KR20 estimate a neutron star mass of $M_{NS}=2.24\pm 0.09$ M using a companion surface optical model that contains a magnetic pole hot spot. Such a massive neutron star would constrain properties of the dense matter equation of state (Lattimer & Prakash, 2007; Steiner et al., 2013; Brandes et al., 2023). Previous modeling is sensitive to the poorly constrained companion hot spot, adding uncertainty to the binary inclination $i$ and, by extension, $M_{NS}$ estimates. Unusually, the previous best-fit companion model has sub-solar metallicity with $\log Z=-1$. Also, older X-ray observations, which suggested widely separated light curve peaks and a flat IBS, relied on an XMM-Newton exposure with strong background flaring and only two binary orbits. Thus, we conduct more detailed optical and X-ray study to better understand the robustness and systematics of the companion heating model and probe the IBS physics.
@@ 2.1 X-rays
We convert the PN and MOS light curves from count rate and count rate error to flux in each detector. We combine the data and distribute it into 26 phase bins. We compute the average flux weighted by the inverse square of the errors in each bin. Flux errors are estimated by the standard error of a variance-weighted sample. Our combined 0.5-10 keV X-ray light curve is shown in the top panel of Fig. 1.
Figure 1: Binned X-ray and optical orbital light curves of J2215. More-finely binned light curves are shown by faded markers. (Top) The combined PN and MOS 0.5-10 keV X-ray light curve. We identify four distinct regions for which we separately fit spectra: two peaks, P1 and P2 corresponding to the two phase intervals in which the flow of the IBS shock is tangent to our line of sight, a bridge (B) between the two peaks, and an off region around the companion inferior conjunction where the X-ray emission comes principally from the IBS nose. (Bottom) The OM U filter light curve for J2215 from the new XMM-Newton observations.
@@ 2.2 Optical
Contemporaneous OM U filter data was taken in both image and fast mode. The fast mode ${}^{\prime\prime}\times 10^{\prime\prime}$ images were obtained from a series of exposures over each observation. We use the SAS tool omfchain to obtain U filter light curves. For each exposure, we extract the source from a circular aperture with a radius of 3 detector pixels in the fast mode image and determine the background from the average count rate of the corresponding image mode image with all point sources removed. The extracted count rates are subsequently corrected by the fraction of the point-spread function (PSF) contained in the aperture. After processing, the OM exposure time is 85 ks. We convert the OM U data to flux using the published OM U calibration for A-type stars. We place the data into flux bins by computing the sample mean and standard error. We also use the Fermi-LAT third gamma-ray pulsar catalog ephemeris for the orbital parameters. The OM U light curve is shown in the bottom panel of Fig. 1.
@@ 3 Spectral Analysis
We divide our X-ray light curve into four phase intervals as labeled in Fig. 1 and perform spectral analyses of these phase bins. We identify Peak 1 (P1; $0.580.44<\Phi<$ 0.58) and Peak 2 (P2; $1.070.87<\Phi<$ 1.07) as the two characteristic IBS X-ray peaks (Romani & Sanchez, 2016; Kandel et al., 2019). We demarcate a bridge region (B; $0.870.58<\Phi<$ 0.87) between P1 and P2 and an off region ($0.440.07<\Phi<$ 0.44) corresponding to the companion at inferior conjunction. Here non-IBS fluxes (e.g. from the magnetosphere or a pulsar wind nebula) may dominate, as only weakly beamed X-rays from the IBS nose are expected. We use the CIAO tool Sherpa (Freeman et al., 2001; Doe et al., 2007) to fit non-thermal power-law spectral models to the XMM data with spectral indices $\Gamma_{P1}$, $\Gamma_{P2}$, and $\Gamma_{B}$. We include an additional phase-independent power-law component throughout the orbit with power-law index $\Gamma_{0}$ to account for phase-independent emission. The spectral model in counts s-1 cm-2 keV-1 in each phase region is
@@
where $N_{H}$ is the equivalent hydrogen column, $\sigma(E)$ is the photo-electric cross section, $K_{r}$ and $K_{0}$ are power-law normalization factors, and $r=\{P1,P2,B,\textit{Off}\}$ demarcates the orbital phase region. For our absorption model, we employ the photoelectric cross sections of Balucinska-Church & McCammon (1992) and the elemental abundances of Anders & Grevesse (1989). In the off phase, $K_{r}=0$. The results of our fit are shown in Table 1. Because the spectral indices of P1, P2, and B are not significantly different, we also give the average spectral index across the IBS $\Gamma_{IBS}$. Fig. 2 shows the spectra across the IBS as well as in the off phases as observed by PN.
Figure 2: The observed PN X-ray spectra with the equivalent hydrogen column $N_{H}$ fixed at $0.09\times 10^{22}$ cm-2. The plotted points are the spectra in the IBS phase region ($1.070.44<\Phi<$ 1.07; including P1, P2, and B) as well as in the off region ($0.440.07<\Phi<$ 0.44). The solid line shows the full model in the IBS phases $f_{tot}$ and the two dashed lines show the absorbed power-law models for the IBS component $f_{IBS}$ with power-law index $\Gamma_{IBS}$ and the phase-independent component $f_{Off}$ with power-law index $\Gamma_{0}$. Note that in the off region, the only model component is the absorbed power-law $f_{Off}$.
@@ 4.1 Optical Models
Previous optical modeling of J2215 has been conducted by Linares et al. (2018) and KR20 using g’r’i’ data from two observing nights at the William Herschel Telescope (WHT). These light curve analyses show strong heating asymmetries on the companion star, suggesting the presence of a hot spot, which could plausibly be a magnetic pole (Sanchez & Romani, 2017). KR20 include a hot spot at the coordinates $(\theta_{HS},\phi_{HS})$ defined from the nose on the stellar surface. At that location, the stellar temperature increases by a factor $(1+A_{HS})$ from $T\approx 5600$ K to $T\gtrsim 8000$ K with a Gaussian profile over radius $r_{HS}$. Their best-fit model gives $i=68.9^{\circ}$ and consequently a very high $M_{NS}=2.24$ M. These results, however, are sensitive to the hot spot parameters. Additionally, they were obtained using a simplified gravity darkening model, which does not account for the gravity darkening of the reprocessed pulsar heating luminosity (Claret & Bloemen, 2011; Kandel & Romani, 2023).
We refit the WHT g’r’i’ data along with the new OM U data using the ICARUS light curve modeling code (Breton et al., 2013) including the extensions described in KR20 and the revised gravity darkening law described in Kandel & Romani (2023). The model is very similar to the best-fit hot spot model of KR20; we augment a direct pulsar heating model with a surface hot spot and add a flat spectrum veiling flux $f_{2}$ to the WHT night 2 data. We adopt an extinction $A_{V}=0.4$ in these fits, as this matches the fit $N_{H}$ (Table 1) and is consistent with estimates from current 3-D dust maps (Green et al., 2018). We perform Monte Carlo sampling of the parameters using the library PyMultinest (Feroz & Hobson, 2008; Feroz et al., 2009; Buchner et al., 2014; Feroz et al., 2019), which uses nested sampling to search the parameter space. To predict the flux in each photometric band, we use BT-Settl model atmospheres (Allard, 2014) available from the Spanish Virtual Observatory (http://svo2.cab.inta-csic.es). These model atmosphere spectra are available with a range of metallicities. We have explored models with $\log Z=-1.5$ to $+0.5$; a few models are also available with varying degrees of $\alpha$ process abundance enhancement.
@@
Our fits prefer calibration offsets $\delta m_{r}\gtrsim 0.03$ to both nights in the WHT r’ band and $\delta m_{U}\approx 0.5$ to the OM U data. We have attempted to find the origin of this large U offset. As noted in Sec. 2.2, we obtain the OM U AB magnitudes by converting from count rate to flux with the SAS-recommended U conversion factor of $1.7\times 10^{-16}$ erg cm-2 Å-1 count-1 for A stars. This spectral type is appropriate for the companion near optical maximum. Using archival XMM OM U exposures in fields covered by the Sloan Digital Sky Survey (SDSS), where catalog $u$ magnitudes are available, we have also investigated the calibration manually. We extracted the count rates in 3-pixel apertures centered on bright stars in two SDSS fields, subtracted the average backgrounds in the images, and scaled the stars’ count rates by the calibration aperture correction factor given by SAS. From these data we find a conversion to SDSS u of $2.7\pm 0.2\times 10^{-16}$ erg cm-2 Å-1 count-1, which would suggest that the U fluxes are brighter by a factor of $\sim 1.6$ (or offset by $\sim 0.5$ mag) with respect to the our XMM calibrated flux, comparable to the fit offset.
@@
We have also examined the calibrated 2014 Keck LRIS spectra of Romani et al. (2015), whose g’r’i’ colors are consistent with the 2014 WHT photometry. We find the Keck fluxes at wavelengths below $4000$ are consistent with the family of BT-Settl model atmosphere spectra (Allard, 2014) at the day side temperature of the star. A change to the thermal companion flux (unlike the IBS non-thermal flux) seems unlikely and would require a concomitant change in the heating power. Our X-ray flux is consistent with past values and there is no significant change to the (dominant) Fermi-LAT gamma-ray heating. Consequently, we consider a U calibration offset the most likely explanation. While fitting for $\delta m_{U}$ slightly decreases the predictive power of our model, the OM U light curve shape still probes the hot spot geometry.
When fitting the g’r’i’ light curve data without including U, we find that the best-fit parameters for $i$ and the hot spot are, unusually, sensitive to the metallicity of the companion stellar atmosphere. This happens for J2215 because the location of the hot spot allows it to mimic a substantial portion of the companion nose heating under the corrected gravity darkening law. With only g’r’i’ colors, the spot location and temperature are not independently constrained, so there is substantial covariance with $i$ in the fits. Consequently, the subtle metallicity-dependent color differences shift the best-fit $i$ and hot spot parameters. For each atmosphere model, we show the $\Delta\chi^{2}$ from the best fitting metallicity model as well as the estimated inclination in Fig. 3. This hot spot metallicity sensitivity seems to be a peculiarity of J2215’s parameters; we do not see such sensitivity in our fits of other objects (Draghis et al., 2019). Nevertheless, this remains a possible rare systematic that should be checked in spider model fitting.
The enhancement of $\alpha$-process elements increases as $\log Z$ decreases in the BT-Settl models. For $\log Z\geq 0$, all models have $\alpha=0$, while for $\log Z\leq-1$ the bulk of the models have $\alpha=0.4$. At $\log Z=-0.5$ most models have $\alpha=0.2$, but there are a subset at $\alpha=0$ that span fewer temperatures and surface gravity values. With the WHT g’r’i’ data, we find that super-solar metallicities are unacceptable while all $\log Z\leq 0$ give comparable $\chi^{2}$ (see Fig. 3). Additional color information can break the degeneracy between atmosphere models. Thus adding XMM U data helps resolve this ambiguity. The best-fit models with the standard $\alpha$ enhancements select slightly sub-solar metallicity with a minimum $\chi^{2}$/DoF=1.35 for $(\log Z,\alpha)=(-1,0.4)$. $(\log Z,\alpha)=(-0.5,0.2)$ has an increased $\chi^{2}$; however, the best fit model improves with the limited $(\log Z,\alpha)=(-0.5,0)$ grid. The $(\log Z,\alpha)=(-0.5,0)$ atmosphere represents the global best fit with $\Delta\chi^{2}=-1$ improvement over $(\log Z,\alpha)=(-1,0.4)$. Inspecting the synthesized colors, we note that the $\alpha$ changes between $(-0.5,0)$ and $(-0.5,0.2)$ introduce small $\sim 5\%$ differences in the emergent flux, especially in the bluest bands for $T<7000$ K. Unfortunately, the online spectra have insufficient $\alpha$ coverage to fully explore this subtle effect.
The optical spectra of Romani et al. (2015) also probe the metallicity. In comparing Ca, Mg and Fe absorption line equivalent widths at maximum with BT-Settl model spectra, we see a best match to $\log Z=0$, similar to the spider companions examined in Draghis et al. (2019). The discrepancy might be related to some imprecision in the colors synthesized from the spectra or in the optical broad band flux calibration. In Table 2, we show the estimated parameters of the Ug’r’i’ models using the $(\log Z,\alpha)=(-1,0.4)$, $(-0.5,0)$, and $(0,0)$ atmospheres. Fig. 4 shows the best-fit light curve for the $(\log Z,\alpha)=(-0.5,0)$ model.
Figure 3: Sensitivity of the light curve fits to $\log Z$. The top panel shows the best-fit inclination, while the lower panel shows the $\Delta\chi^{2}$ from minimum. Fits of gri alone are shown in red while fits of Ugri are shown in blue. For gri all $\log Z\leq 0$ metallicities are acceptable with near-identical $\chi^{2}$, leaving large $i$ (and mass) uncertainty. Including XMM U breaks this degeneracy, preferring $\log Z=-1$ (the value used in KR20). The left points of the $\log Z=-0.5$ pair are for a lower $\alpha=0$ element abundance; these represent the global minimum values found in the fits. See the text for details on $\alpha$ abundance.
Figure 4: The best-fit companion optical light curve model plotted against the gri data from WHT and the OM U data shifted by the estimated calibration offsets along with the residuals. The companion has $(\log Z,\alpha)=(-0.5,0)$. Phase $0\leq\Phi\leq 1$ shows WHT night 1 while $1\leq\Phi\leq 2$ shows WHT night 2.
The larger values of $i$ now preferred imply a lower neutron star mass than found by KR20. To determine this mass (also shown in Table 2), we use the KR20 radial velocity measurement $K_{c}=429.8\pm 3.9$ km/s. Including $1\sigma$ statistical uncertainties, our mass estimates range from $M_{NS}=1.85-2.3$ M, increasing with decreased metallicity. Although deprecated by the photometry, the spectroscopically preferred $\log Z=0$ gives $1.91\pm 0.06M_{\odot}$ while the photometrically acceptable $\log Z=-1$ fit gives $M_{NS}=2.18\pm 0.10$ M. These are both well less than the ($\log Z=-1$) $M_{NS}=2.24\pm 0.09$ M estimate of KR20.
Parameters -1.0/0.4 $-0.5/0$ 0/0 KR20
@@
$\delta m_{r}^{\dagger}$ $0.027(1)$ $0.035(1)$ 0.044(1) $0.031(1)$
$\theta_{HS}$ (deg) $339.3^{+4.8}_{-4.8}$ $322.6^{+5.9}_{-5.8}$ $302.2^{+3.9}_{-3.8}$ $324.0^{+10.1}_{-8.7}$
[…]
$A_{HS}$ $0.31^{+0.10}_{-0.07}$ $0.45^{+0.18}_{-0.13}$ $0.90^{+0.33}_{-0.29}$ $0.6^{+0.4}_{-0.2}$
[…]
$\delta m_{U}^{\dagger}$ 0.472(13) 0.477(13) 0.352(14)
$\chi^{2}$/DoF 351/260 349/260 388/260 297/232
Table 2: Parameter results for optical modeling of WHT+OM J2215 data including the 68% confidence intervals. We show the best-fitting results with metallicities $(\log Z,\alpha)=(-1,0.4)$, $(-0.5,0)$, and $(0,0)$. $f_{c}$ is the Roche lobe filling factor, $L_{H,34}$ is the pulsar heating luminosity for an equatorially concentrated heating flux in units of $10^{34}\,{\rm erg\,s^{-1}}$, and $d$ is the source distance. The hot spot coordinates $(\theta_{HS},\phi_{HS})$ are defined so that $(0,0)$ refers to the sub-pulsar point at the companion nose. The right column shows the fit results of KR20 which used $\log Z=-1$, assumed a different gravity darkening model, and did not have U data. $\delta$ are band offsets with last digit errors. $\delta m_{U}$ is from the standard SAS flux calibration; however, we see evidence for a systematic calibration shift of $\sim$0.5 mag. These fits assume $A_{V}=0.40$, consistent with KR20, our X-ray fits and 3-D dust maps. We have conducted limited (expensive) tests with free $A_{V}$ in fits, which shift the solutions to slightly larger $A_{V}$ with higher $L_{H}$ and $T_{N}$ to correct the colors; other parameters remain within $1\sigma$ of the fixed $A_{V}$ fit values.